Polaski Company manufactures and sells a single product called a Ret. Operating at capacity, the company can produce and sell1. Financial advantage 2. Financial advantage 3. Financial (disadvantage)

HIGHLIGHT ANSWERS!! HIGHLIGHT ANSWERS!! HIGHLIGHT ANSWERS!! HIGHLIGHT ANSWERS!!
Answer #1

Part 1

Financial advantage (disadvantage) = sales – variable costs – special machine costs = (8000*51*(1-16%))-(8000*(20+8+3+(4*(1-75%))))-16000 = 70720

Therefore,

Financial advantage

$70720

Part 2

Financial advantage (disadvantage) = reimbursement for costs of production + fixed fee – less incremental costs = (8000*(20+8+3+5))+(1.40*8000)-(8000*(20+8+3)) = 51200

Therefore,

Financial advantage

$51200

Part 3

Financial advantage (disadvantage) = from the US army – from regular channels – variable selling expense = (8000*(20+8+3+5))-(8000*51)-(8000*4) = $152000

Therefore,

Financial disadvantage

$(152000)